OpenStudy (anonymous):

Find the smallest value of a so that the function y=e^(2x) *x^2 * a*e^(2x)+3 would be growing no matter what the x value is.

OpenStudy (anonymous):

The answer given is 1.

OpenStudy (anonymous):

Sorry, I made a mistake. The function is y=e^(2x) *x^2+a*e^(2x)+3

OpenStudy (experimentx):

what does "would be growing no matter what the x value is" mean??

OpenStudy (anonymous):

The function gets bigger not smaller.

OpenStudy (experimentx):

a>0

OpenStudy (anonymous):

Oh, alright. Thanks!

OpenStudy (experimentx):

seems to be growing for a=0.001

OpenStudy (experimentx):

wait thats not correct

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=+y%3De%5E%282x%29+%28x%5E2+-+25%29+%2B3 seems to me that the function is always growing after certain interval of time

OpenStudy (experimentx):

I think a=1 is the right answer ...

OpenStudy (anonymous):

But how do you get it?

OpenStudy (experimentx):

for a=0, we have y=e^(2x) (x^2) +3 and this (x^2) is less than 1 between 0 to 1 and is decreasing the function.

OpenStudy (anonymous):

Oh, alright.

OpenStudy (experimentx):

@ErkoT you still there ??? still I am quite incorrect

OpenStudy (experimentx):

y=e^(2x) (x^2) +3 increases between 0 to 1 it decreases on -1 to 0

OpenStudy (experimentx):

ah ... i found the best way, take derivative of e^(2x) (x^2+a) and it must be > 0, and find the value of a as x->-0